OpenStudy (anonymous):

A 14 gram sample of a substance that's used to preserve fruit and vegetables has a k-value of 0.1092. Find the substance's half-life, in days. Round your answer to the nearest tenth.

OpenStudy (ribhu):

half life period = ln2 / k = 0.693/0.1092 = 6.35 days

OpenStudy (anonymous):

That was not the answer, but thanks anyway.

OpenStudy (anonymous):

6.3 is the answer